# Information Limits of Joint Community Detection and Finite Group Synchronization

### v1

@@ 5.1 Preliminaries
To proceed, recall the definition of the two subsets sets $S_{12}$ and $S_{21}$ in Section 5.1 which includes the nodes with different labels between $\bm{\kappa}$ and $\bm{\kappa}^{*}$. Let us denote $\Delta:=|S_{12}|=|S_{21}|$ as their sizes, then for any wrong hypothesis $\bm{\kappa}$ with $\text{dist}_{\mathrm{c}}(\bm{\kappa},\bm{\kappa}^{*})>0$, it satisfies $11\leq\Delta\leq\frac{n}{2}-$ 1 (when $\Delta=\frac{n}{2}$, exact recovery can be achieved by swapping the labels of $S_{1}$ and $S_{2}$). In this way, for a fixed $\Delta$, we define
@@
as the set of all such hypotheses, and there are $\binom{n/2}{\Delta}^{2}$ number of unique $\bm{\kappa}$, i.e., $|\mathcal{H}_{\bm{\kappa},\Delta}|=\binom{n/2}{\Delta}^{2}.$ Furthermore, let $\bm{\kappa}_{\Delta}\in\mathcal{H}_{\bm{\kappa},\Delta}$ be any hypothesis in $\mathcal{H}_{\bm{\kappa},\Delta}$, we denote
$\mathcal{H}_{\bm{x},\Delta}:=\{\bm{x}=\{\bm{\kappa},\bm{g}\}:\bm{\kappa}=\bm{\kappa}_{\Delta}\},$
as the set of hypotheses $\bm{x}$ with a fixed $\bm{\kappa}$ and different group elements $\bm{g}$. Given the above, the error probability in (33) can be bounded as
@@
=$\overset{(c)}{=}\sum_{\Delta=1}^{n/2-1}\binom{n/2}{\Delta}^{2}\underbrace{\mathcal{P}_{\bm{x}^{*}}\left(\exists\bm{x}\in\mathcal{H}_{\bm{\kappa},\Delta}:\mathcal{P}_{\bm{x}}(\bm{y})\geq\mathcal{P}_{\bm{x}^{*}}(\bm{y})\right)}_{=:\mathcal{P}_{\Delta}}$
@@
Here, $(a)$ holds by applying the union bound over different $\Delta$; $(b)$ comes by further applying this over the set $\mathcal{H}_{\bm{\kappa},\Delta}$; $(c)$ stems from the fact that any hypothesis $\bm{\kappa}\in\mathcal{H}_{\bm{\kappa},\Delta}$ should contribute the same amount of error probability under a fixed $\Delta$; $(d)$ uses the symmetry between $\Delta$ and $\frac{n}{2}-\Delta$ by swapping the labels in $\bm{\kappa}$.
Step 2: Bound on each $\mathcal{P}_{\Delta}$
@@
$\mathcal{P}_{\Delta}$ $=\mathcal{P}_{\bm{x}^{*}}\left(\exists\bm{x}\in\mathcal{H}_{\bm{\kappa},\Delta}:\frac{\mathcal{P}_{\bm{x}}(\bm{y})}{\mathcal{P}_{\bm{x}^{*}}(\bm{y})}\geq 1\right)$
$\leq\mathcal{P}_{\bm{x}^{*}}\left(\exists\bm{x}\in\mathcal{H}_{\bm{\kappa},\Delta}:\left(\frac{Ma}{b}\right)^{r-r^{*}}\cdot\frac{\mathbbm{1}(g_{ij}=g_{i}g_{j}^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}})}{\mathbbm{1}(g_{ij}=g_{i}^{*}(g_{j}^{*})^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}^{*})}>1\right)$
$=\mathcal{P}_{\bm{x}^{*}}\left(\exists\bm{x}\in\mathcal{H}_{\bm{\kappa},\Delta}:\left(\frac{Ma}{b}\right)^{r-r^{*}}\cdot\mathbbm{1}(g_{ij}=g_{i}g_{j}^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}})>1\right)$ (35)
@@
$\mathcal{P}_{\bm{x}^{*}}\left(g_{ij}=g_{i}^{*}(g_{j}^{*})^{-1};\;\forall(i,j)\in\mathcal{E}^{*}_{\text{inner}}\right)=1.$
@@
To move forward, for any wrong hypothesis $\bm{x}$, let us define the following two events in the probability space:
$E_{1}$ $:\quad\left(\frac{Ma}{b}\right)^{r-r^{*}}>1,$ (36)
$E_{\bm{x},2}$ $:\quad g_{ij}=g_{i}g_{j}^{-1},\;\forall(i,j)\in\mathcal{E}_{\text{inner}}.$
which are the necessary and sufficient conditions for $\bm{x}$ with a likelihood higher than $\bm{x}^{*}$. Notably, for different hypothesis $\bm{x}\in\mathcal{H}_{\bm{\kappa},\Delta}$ with the fixed community assignment $\bm{\kappa}=\bm{\kappa}_{\Delta}$, the event $E_{1}$ does not depend on $\bm{x}$ but is invariant in the sample space, therefore we omit the subscript $\bm{x}$. In this way, (35) can be written as
@@
$\mathcal{P}_{\Delta}$ $\leq\mathcal{P}_{\bm{x}^{*}}\left(\bigcup\limits_{\bm{x}\in\mathcal{H}_{\bm{\kappa},\Delta}}\left(E_{1}\bigcap E_{\bm{x},2}\right)\right)=\mathcal{P}_{\bm{x}^{*}}\left(E_{1}\bigcap\underbrace{\left(\bigcup\limits_{\bm{x}\in\mathcal{H}_{\bm{\kappa},\Delta}}E_{\bm{x},2}\right)}_{=:E_{2}}\right)$
$=\mathcal{P}_{\bm{x}^{*}}\left(E_{1}\bigcap E_{2}\right)$ (37)
where we denote $E_{2}$ as the union over $\bm{x}\in\mathcal{H}_{\bm{\kappa},\Delta}$. To proceed, recall the definitions of $r$ and $r^{*}$ in (30), we further define $r=r_{1}+r_{2}$ where
$r_{1}:=|\mathcal{E}(S_{11},S_{12})|,\quad r_{2}:=|\mathcal{E}(S_{21},S_{22})|.$
Under the ground truth $\bm{x}^{*}$, the three $r_{1}$, $r_{2}$, and $r^{*}$ are independent binomial random variables since they involve mutually disjoint sets of edges. Therefore, by conditioning on $r_{1}$, $r_{2}$, and $r^{*}$, (37) satisfies
@@
$\mathcal{P}_{\Delta}$ $\leq\sum_{r_{1}}\sum_{r_{2}}\sum_{r^{*}}\mathcal{P}_{\bm{x}^{*}}(r_{1})\mathcal{P}_{\bm{x}^{*}}(r_{2})\mathcal{P}_{\bm{x}^{*}}(r^{*})\mathcal{P}_{\bm{x}^{*}}\left(E_{1}\bigcap E_{2}\mid r_{1},r_{2},r^{*}\right)$
$=\sum_{r_{1}}\sum_{r_{2}}\sum_{r^{*}}\mathcal{P}_{\bm{x}^{*}}(r_{1})\mathcal{P}_{\bm{x}^{*}}(r_{2})\mathcal{P}_{\bm{x}^{*}}(r^{*})\mathcal{P}_{\bm{x}^{*}}\left(\left(\left(\frac{Ma}{b}\right)^{r_{1}+r_{2}-r^{*}}>1\right)\bigcap E_{2}\mid r_{1},r_{2},r^{*}\right)$
$=\sum_{r_{1}}\sum_{r_{2}}\sum_{r^{*}}\mathcal{P}_{\bm{x}^{*}}(r_{1})\mathcal{P}_{\bm{x}^{*}}(r_{2})\mathcal{P}_{\bm{x}^{*}}(r^{*})\cdot\mathbbm{1}\left(\left(\frac{Ma}{b}\right)^{r_{1}+r_{2}-r^{*}}>1\right)\cdot\mathcal{P}_{\bm{x}^{*}}\left(E_{2}\mid r_{1},r_{2},r^{*}\right)$ (38)
@@
where the last step comes from the fact that given $r_{1}$, $r_{2}$, and $r^{*}$, the probability $\mathcal{P}_{\bm{x}^{*}}(E_{1}\mid r_{1},r_{2},r^{*})$ is either 0 or 1. Therefore we can represent it as an indicator function. Furthermore, by introducing an extra factor $\alpha>0$, one can see that it satisfies
$\mathbbm{1}\left(\left(\frac{Ma}{b}\right)^{r_{1}+r_{2}-r^{*}}>1\right)$ $=\mathbbm{1}\left(\left(\frac{Ma}{b}\right)^{\alpha(r_{1}+r_{2}-r^{*})}>1\right)$
$\leq\left(\frac{Ma}{b}\right)^{\alpha(r_{1}+r_{2}-r^{*})}.$
@@
$\mathcal{P}_{\Delta}\leq\min_{\alpha>0}\left(\sum_{r_{1}}\sum_{r_{2}}\sum_{r^{*}}\mathcal{P}_{\bm{x}^{*}}(r_{1})\mathcal{P}_{\bm{x}^{*}}(r_{2})\mathcal{P}_{\bm{x}^{*}}(r^{*})\cdot\left(\frac{Ma}{b}\right)^{\alpha(r_{1}+r_{2}-r^{*})}\cdot\mathcal{P}_{\bm{x}^{*}}\left(E_{2}\mid r_{1},r_{2},r^{*}\right)\right).$ (39)
Notably, $\alpha$ is significant for obtaining the sharp result since we can minimize over all $\alpha>0$.
Step 3: Cycle analysis
Now, we need to upper bound $\mathcal{P}_{\bm{x}^{*}}\left(E_{2}\mid r_{1},r_{2},r^{*}\right)$. Importantly, a naive way by applying the union bound over all possible hypotheses in $\mathcal{H}_{\bm{x},\Delta}$ such that
$\mathcal{P}_{\bm{x}^{*}}\left(E_{2}\mid r_{1},r_{2},r^{*}\right)$ $\leq\sum_{\bm{x}\in\mathcal{H}_{\bm{\kappa},\Delta}}\mathcal{P}_{\bm{x}^{*}}\left(E_{\bm{x},2}\mid r_{1},r_{2},r^{*}\right)$
fails to give a sharp result, since $|\mathcal{H}_{\bm{\kappa},\Delta}|=M^{n}$ where each node has $M$ choices of its group elements, then the summation over $\mathcal{H}_{\bm{\kappa},\Delta}$ explodes as $n$ is large. Instead, we have to bound $\mathcal{P}_{\bm{x}^{*}}\left(E_{2}\mid r_{1},r_{2},r^{*}\right)$ directly as a whole. To this end, let us restate the event $E_{2}$ defined in (37) as
$E_{2}:\quad\exists\bm{x}\in\mathcal{H}_{\bm{\kappa},\Delta}:\;g_{ij}=g_{i}g_{j}^{-1},\;\forall(i,j)\in\mathcal{E}_{\text{inner}}.$ (40)
Here, recall $\mathcal{E}_{\text{inner}}$ denotes the set of edges within the two communities $S_{1}$ and $S_{2}$, i.e. $\mathcal{E}_{\text{inner}}=\mathcal{E}(S_{1})\cup\mathcal{E}(S_{2})$, then we can consider $S_{1}$ and $S_{2}$ separately by further defining the following two events:
$E_{2,S_{1}}:$ $\quad\exists\bm{x}\in\mathcal{H}_{\bm{\kappa},\Delta}:g_{ij}=g_{i}g_{j}^{-1},\;\forall(i,j)\in\mathcal{E}(S_{1}),$ (41)
$E_{2,S_{2}}:$ $\quad\exists\bm{x}\in\mathcal{H}_{\bm{\kappa},\Delta}:g_{ij}=g_{i}g_{j}^{-1},\;\forall(i,j)\in\mathcal{E}(S_{2}),$

### v2

@@ 6.1 Step 1: A union bound
To proceed, recall the definition of the two subsets $S_{12}$ and $S_{21}$ in (28), which denotes the set of nodes assigned with different labels between $\bm{\kappa}$ and $\bm{\kappa}^{*}$. Let us further define
$\Delta:=|S_{12}|=|S_{21}|$ (40)
as their sizes, then for any wrong hypothesis $\bm{\kappa}$ with $\text{dist}_{\mathrm{c}}(\bm{\kappa},\bm{\kappa}^{*})>0$, it must satisfy $11\leq\Delta\leq\frac{n}{2}-$ 1. When $\Delta=\frac{n}{2}$, exact recovery can be achieved by swapping the labels of $S_{1}$ and $S_{2}$. In this way, for a fixed $\Delta$, we define
@@
as the set of all satisfying hypotheses (of the cluster memberships), which includes $\binom{n/2}{\Delta}^{2}$ number of unique $\bm{\kappa}$, i.e.,
$|\mathcal{H}_{\bm{\kappa},\Delta}|=\binom{n/2}{\Delta}^{2}.$
Furthermore, let $\bm{\kappa}_{\Delta}\in\mathcal{H}_{\bm{\kappa},\Delta}$ be any fixed hypothesis in $\mathcal{H}_{\bm{\kappa},\Delta}$, the error probability in (39) can be upper bounded as
@@
=$\color[rgb]{0,0,0}{\overset{(c)}{=}\sum_{\Delta=1}^{n/2-1}\binom{n/2}{\Delta}^{2}\underbrace{\mathcal{P}_{\bm{x}^{*}}\left(\exists\bm{x}=\{\bm{\kappa}_{\Delta},\bm{g}\}:\mathcal{P}_{\bm{x}}(\bm{y})\geq\mathcal{P}_{\bm{x}^{*}}(\bm{y})\right)}_{=:\mathcal{P}_{\Delta}}}$
@@
Here, $(a)$ holds by applying the union bound over different $\Delta$; $(b)$ comes by further applying this over the set $\mathcal{H}_{\bm{\kappa},\Delta}$; $(c)$ stems from the fact that any hypothesis $\bm{\kappa}\in\mathcal{H}_{\bm{\kappa},\Delta}$ should contribute the same amount of error probability under a fixed $\Delta$ and therefore we focus on $\bm{\kappa}=\bm{\kappa}_{\Delta}$ as one of them; $(d)$ uses the symmetry between $\Delta$ and $\frac{n}{2}-\Delta$ by swapping the labels in $\bm{\kappa}$.
6.2 Step 2: Bound on each $\mathcal{P}_{\Delta}$
@@
$\mathcal{P}_{\Delta}$ $=\mathcal{P}_{\bm{x}^{*}}\left(\exists\bm{x}=\{\bm{\kappa}_{\Delta},\bm{g}\}:\frac{\mathcal{P}_{\bm{x}}(\bm{y})}{\mathcal{P}_{\bm{x}^{*}}(\bm{y})}\geq 1\right)$
$=\mathcal{P}_{\bm{x}^{*}}\left(\exists\bm{x}=\{\bm{\kappa}_{\Delta},\bm{g}\}:\left(\frac{Ma(1-q)}{b(1-p)}\right)^{r-r^{*}}\cdot\frac{\mathbbm{1}(g_{ij}=g_{i}g_{j}^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}})}{\mathbbm{1}(g_{ij}=g_{i}^{*}(g_{j}^{*})^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}^{*})}\geq 1\right)$
$=\mathcal{P}_{\bm{x}^{*}}\left(\exists\bm{x}=\{\bm{\kappa}_{\Delta},\bm{g}\}:\left(\frac{Ma(1-q)}{b(1-p)}\right)^{r-r^{*}}\cdot\mathbbm{1}(g_{ij}=g_{i}g_{j}^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}})\geq 1\right),$ (42)
@@
$\mathcal{P}_{\bm{x}^{*}}\left(g_{ij}=g_{i}^{*}(g_{j}^{*})^{-1};\;\forall(i,j)\in\mathcal{E}^{*}_{\text{inner}}\right)=1,$
based on our model setting that any group transformation $g_{ij}$ is noiseless when $(i,j)\in\mathcal{E}^{*}_{\text{inner}}$.
@@
To move forward, (42) motivates us to define the following two events in the probability space:
$E_{1,\bm{x}}$ $:\quad\left(\frac{Ma(1-q)}{b(1-p)}\right)^{r-r^{*}}\geq 1,\quad\text{for }\bm{x},$ (43)
$E_{2,\bm{x}}$ $:\quad g_{ij}=g_{i}g_{j}^{-1},\;\forall(i,j)\in\mathcal{E}_{\text{inner}},\quad\text{for }\bm{x}.$
In this way, (42) essentially measures the probability of having any $\bm{x}=\{\bm{\kappa}_{\Delta},\bm{g}\}$ that makes the joint event $E_{1,\bm{x}}\cap E_{2,\bm{x}}$ happen. Therefore, (42) can be rewritten as
@@
$\mathcal{P}_{\Delta}=\mathcal{P}_{\bm{x}^{*}}\left(\bigcup\limits_{\bm{x}:\bm{\kappa}=\bm{\kappa}_{\Delta}}\left(E_{1,\bm{x}}\bigcap E_{2,\bm{x}}\right)\right).$ (44)
Furthermore, notice that for any $\bm{x}=\{\bm{\kappa}_{\Delta},\bm{g}\}$ where the cluster membership $\bm{\kappa}=\bm{\kappa}_{\Delta}$ is fixed, the event $E_{1,\bm{x}}$ is actually invariant to the choice of $\bm{g}$ as it only involves edge connections (recall $r$ and $r^{*}$ defined in (36)). Therefore, we can define
$E_{1}:\quad\left(\frac{Ma(1-q)}{b(1-p)}\right)^{r-r^{*}}\geq 1,\quad\text{for }\bm{\kappa}=\bm{\kappa}_{\Delta},$
that does not depend on $\bm{x}$ and further express $\mathcal{P}_{\Delta}$ as
@@
$\mathcal{P}_{\Delta}$ $=\mathcal{P}_{\bm{x}^{*}}\left(\bigcup\limits_{\bm{x}:\bm{\kappa}=\bm{\kappa}_{\Delta}}\left(E_{1,\bm{x}}\bigcap E_{2,\bm{x}}\right)\right)=\mathcal{P}_{\bm{x}^{*}}\left(\bigcup\limits_{\bm{x}:\bm{\kappa}=\bm{\kappa}_{\Delta}}\left(E_{1}\bigcap E_{2,\bm{x}}\right)\right)=\mathcal{P}_{\bm{x}^{*}}\left(E_{1}\bigcap\underbrace{\left(\bigcup\limits_{\bm{x}:\bm{\kappa}=\bm{\kappa}_{\Delta}}E_{2,\bm{x}}\right)}_{=:E_{2}}\right)$
$=\mathcal{P}_{\bm{x}^{*}}\left(E_{1}\bigcap E_{2}\right),$ (45)
where we denote $E_{2}$ as the union of $E_{2,\bm{x}}$.
To proceed, recall the definitions of $r$ and $r^{*}$ in (36), we further split $r$ into $r=r_{1}+r_{2}$ such that
$r_{1}:=|\mathcal{E}(S_{11},S_{12})|,\quad r_{2}:=|\mathcal{E}(S_{21},S_{22})|$ (46)
Under the ground truth $\bm{x}^{*}$, the three random variables $r_{1}$, $r_{2}$, and $r^{*}$ are binomial and independent as they involve mutually disjoint sets of edges (one can refer to Figure 4 for visualization). Therefore, by conditioning on $r_{1}$, $r_{2}$, and $r^{*}$, (45) satisfies
$\mathcal{P}_{\Delta}$ $=\sum_{r_{1}}\sum_{r_{2}}\sum_{r^{*}}\mathcal{P}_{\bm{x}^{*}}(r_{1})\mathcal{P}_{\bm{x}^{*}}(r_{2})\mathcal{P}_{\bm{x}^{*}}(r^{*})\mathcal{P}_{\bm{x}^{*}}\left(E_{1}\bigcap E_{2}\mid r_{1},r_{2},r^{*}\right)$
$=\sum_{r_{1}}\sum_{r_{2}}\sum_{r^{*}}\mathcal{P}_{\bm{x}^{*}}(r_{1})\mathcal{P}_{\bm{x}^{*}}(r_{2})\mathcal{P}_{\bm{x}^{*}}(r^{*})\mathcal{P}_{\bm{x}^{*}}\left(\left(\left(\frac{Ma(1-q)}{b(1-p)}\right)^{r_{1}+r_{2}-r^{*}}>1\right)\bigcap E_{2}\mid r_{1},r_{2},r^{*}\right)$
$=\sum_{r_{1}}\sum_{r_{2}}\sum_{r^{*}}\mathcal{P}_{\bm{x}^{*}}(r_{1})\mathcal{P}_{\bm{x}^{*}}(r_{2})\mathcal{P}_{\bm{x}^{*}}(r^{*})\cdot\mathbbm{1}\left(\left(\frac{Ma(1-q)}{b(1-p)}\right)^{r_{1}+r_{2}-r^{*}}>1\right)\cdot\mathcal{P}_{\bm{x}^{*}}\left(E_{2}\mid r_{1},r_{2},r^{*}\right),$ (47)
@@
where the last step comes from the fact that given $r_{1}$, $r_{2}$, and $r^{*}$, the probability $\mathcal{P}_{\bm{x}^{*}}(E_{1}\mid r_{1},r_{2},r^{*})$ is either 0 or 1. Therefore, we represent it as an indicator function. Furthermore, by introducing an extra factor $\alpha>0$, one can see that it satisfies
$\mathbbm{1}\left(\left(\frac{Ma(1-q)}{b(1-p)}\right)^{r_{1}+r_{2}-r^{*}}>1\right)$ $=\mathbbm{1}\left(\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha(r_{1}+r_{2}-r^{*})}>1\right)$
$\leq\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha(r_{1}+r_{2}-r^{*})}.$
@@
$\mathcal{P}_{\Delta}\leq\min_{\alpha>0}\left(\sum_{r_{1}}\sum_{r_{2}}\sum_{r^{*}}\mathcal{P}_{\bm{x}^{*}}(r_{1})\mathcal{P}_{\bm{x}^{*}}(r_{2})\mathcal{P}_{\bm{x}^{*}}(r^{*})\cdot\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha(r_{1}+r_{2}-r^{*})}\cdot\mathcal{P}_{\bm{x}^{*}}\left(E_{2}\mid r_{1},r_{2},r^{*}\right)\right).$ (48)
Notably, $\alpha$ is crucial for obtaining the sharp result, as we will minimize over all $\alpha>0$ later on.
6.3 Step 3: Cycle consistency analysis
Now, we need to upper bound $\mathcal{P}_{\bm{x}^{*}}\left(E_{2}\mid r_{1},r_{2},r^{*}\right)$. Importantly, a naive way by applying the union bound over all possible hypotheses $\bm{x}=\{\bm{\kappa}_{\Delta},\bm{g}\}$ such as
$\mathcal{P}_{\bm{x}^{*}}\left(E_{2}\mid r_{1},r_{2},r^{*}\right)=\mathcal{P}_{\bm{x}^{*}}\left(\bigcup\limits_{\bm{x}:\bm{\kappa}=\bm{\kappa}_{\Delta}}E_{2,\bm{x}}\mid r_{1},r_{2},r^{*}\right)\leq\sum_{\bm{x}:\bm{\kappa}=\bm{\kappa}_{\Delta}}\mathcal{P}_{\bm{x}^{*}}\left(E_{2,\bm{x}}\mid r_{1},r_{2},r^{*}\right)$
fails to give a sharp result, since the summation over $\bm{x}:\bm{\kappa}=\bm{\kappa}_{\Delta}$ includes $M^{n}$ terms as each node has $M$ choices of its group elements. As a result, the upper bound increases exponentially as $n$ becomes large.
Instead, we have to bound $\mathcal{P}_{\bm{x}^{*}}\left(E_{2}\mid r_{1},r_{2},r^{*}\right)$ directly as a whole. Let us restate the event $E_{2}$ defined in (45) as
$E_{2}:\quad\exists\bm{x}=\{\bm{\kappa}_{\Delta},\bm{g}\}:\;g_{ij}=g_{i}g_{j}^{-1},\;\forall(i,j)\in\mathcal{E}_{\text{inner}}.$ (49)
Here, recall $\mathcal{E}_{\text{inner}}$ denotes the set of edges within the communities $S_{1}$ and $S_{2}$, i.e., $\mathcal{E}_{\text{inner}}=\mathcal{E}(S_{1})\cup\mathcal{E}(S_{2})$, then we can consider $S_{1}$ and $S_{2}$ separately by further defining the following two events:
$E_{2,S_{1}}:$ $\quad\exists\bm{x}=\{\bm{\kappa}_{\Delta},\bm{g}\}:g_{ij}=g_{i}g_{j}^{-1},\;\forall(i,j)\in\mathcal{E}(S_{1}),$ (50)
$E_{2,S_{2}}:$ $\quad\exists\bm{x}=\{\bm{\kappa}_{\Delta},\bm{g}\}:g_{ij}=g_{i}g_{j}^{-1},\;\forall(i,j)\in\mathcal{E}(S_{2}),$